# Uniform subelliptic estimates for degenerating Fokker-Planck equations

### v1

@@ 2. The model operator and the case $\epsilon=1$
is subelliptic since it satisfies the Hörmander condition, see [6] and [8]. By a result of Folland [2], it has a unique fundamental solution that is homogeneous of degree 2-2. In our case there is an elementary formula for $K_{0}(y)$,
$K_{0}(y)=G_{y_{0}}(y^{\prime})G_{\frac{1}{12}y_{0}^{3}}(y^{\prime\prime}),$
@@ 3. The Fokker-Planck equation
Since $A_{2}(0,y^{\prime},y^{\prime\prime})=\partial_{y_{0}}A_{2}(0,y^{\prime},y^{\prime\prime})=0$, we can write $A_{2}(y)=y_{0}^{2}A_{0}(y)$ and $A_{1}(y)=y_{0}A_{0}(y)$, in each case with $A_{0}\in C^{0}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},w}$.
@@ 5. Operator bounds
6. Proof of Theorem 1.1
For simplicity we consider data $u_{0}(x,\xi)\in\mathcal{S}(\mathbb{R}^{2n})$; the a priori estimates will allow the extension of Theorem 1.1 to spaces of functions with $L^{p}$ bounds on a finite number of derivatatives. We use the following two results.
Take $Q$ as in (1.2), and let $u=e^{tQ}u_{0}$. Then
Maximum Principal: $\sup_{t>0}\sup_{x,\xi}u(t,x,\xi)\leq\sup_{x,\xi}u_{0}(x,\xi).$
Mass Conservation: for $t>0$, $\;\int_{\mathbb{R}^{2n}}u(t,x,\xi)\,dx\,d\xi=\int_{\mathbb{R}^{2n}}u_{0}(x,\xi)\,dx\,d\xi.$
Both of these follow as for the standard heat equation; for completeness we include the proofs here. For the maximum principle, it suffices to work on a strip $t\in[0,T]$ for $T<\infty$. Given $c>0$ consider the function $u_{c}=u-ct$, so $(\partial_{t}-Q)u_{c}=-c<0$. At a local maximum of $u_{c}$ over $0\leq t\leq T$ we must have $\nabla_{x}u=0=\nabla_{\xi}u$, and $X_{j}^{2}u\leq 0$ for $j\geq 1$. Additionally $\partial_{t}u\geq 0$ if $t\neq 0$. This leads to a contradiction if $t\neq 0$. Thus $u_{c}$ attains it maximum at $t=0$. Since this holds for all $c>0$, it follows that $u$ attains its maximum at $t=0$.
@@
$\partial_{t}\int u\,dx\,d\xi=\int Qu\,dx\,d\xi.$
The first order terms in $Q$ are divergence free, and the second order terms have constant coefficients, hence we may integrate by parts, using decay of $u$ in $(x,\xi)$, to see that the right hand side vanishes.
@@
$\|e^{tQ}u_{0}\|_{L^{\infty}(\mathbb{R}^{2n})}\leq\|u_{0}\|_{L^{\infty}(\mathbb{R}^{2n})}.$
@@
(6.1) $\|e^{tQ}u_{0}\|_{L^{p}(\mathbb{R}^{2n})}\leq\|u_{0}\|_{L^{p}(\mathbb{R}^{2n})}\quad\text{for all}\;\;1\leq p\leq\infty.$
@@
Given $N$, let $T_{K}$ be a right inverse for $\partial_{t}-Q$ as in (3.5) such that the remainder is of order 3N-3N, and let $g=T_{K}(\delta\otimes u_{0})$. Since $K(v,y)-K_{0,\epsilon}(y)$ is of negative order, and $K_{0,\epsilon}(y_{0},\cdot)\rightarrow\delta(\cdot)$ as $y_{0}\rightarrow 0+$, by (5.1) we have that $\lim_{t\rightarrow 0+}g(t,\cdot)=u_{0}$ in the $L^{p}(\mathbb{R}^{2n})$ norm.
@@
(6.2) $\sup_{t>0}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3m}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}g(t,\cdot)\bigr\|_{L^{p}}\\ \leq C_{N}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3m}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}u_{0}\bigr\|_{L^{p}}.$
Additionally, $(\partial_{t}-Q)g=R_{3N+1}(\delta\otimes u_{0})$ satisfies
(6.3) $\sup_{t>0}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3N}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}(\partial_{t}-Q)g(t,\cdot)\bigr\|_{L^{p}}\leq C_{N}\|u_{0}\|_{L^{p}}.$
Let $h=e^{tQ}u_{0}-g$, so $(\partial_{t}-Q)h=-(\partial_{t}-Q)g$ and $h=0$ for $t\leq 0$. To control $h$, let $K$ be the left parametrix constructed in Theorem 4.2 for $\partial_{t}-Q$ with remainder $R_{3N+1}$ of order 3N-3N, and write
@@
By (6.1) and (6.2) with $m=0$ we have $\|h(t,\cdot)\|_{L^{p}}\leq C\|u_{0}\|_{L^{p}}$ for $t\geq 0$. If $|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq N$, and $w=(t,x,\xi)$, Theorem 5.5 with $m=0$ then yields
$\sup_{t>0}\,\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}\int R_{3N+1}(\epsilon,w,-y)h(\exp_{w}(y))\,dy\,\bigr\|_{L^{p}(dx\,d\xi)}\\ \leq C\|u_{0}\|_{L^{p}},$
@@
$\sup_{t>0}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3N}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}h(t,\cdot)\bigr\|_{L^{p}(dx\,d\xi)}\leq C\|u_{0}\|_{L^{p}}.$
@@ 7. Applications to classical-quantum correspondence
We show here that certain results from [3] and [7] hold in the subelliptic setting considered in this paper. We consider real valued $V$ satisfying (1.1), and complex valued $\ell_{j}$ satisfying (1.3) with non-degenerate $B$, and define $Q$ by (1.2). Our first application builds upon results of [7] concerning closeness in the trace norm of the quantum and classical evolution of certain semi-classical observables. The symbol class $S^{L^{q}}_{\rho}$ denotes symbols $a(h,x,\xi)$ such that, for all multi-indices $\alpha$, there is $C_{\alpha}$ so that
This space of symbols is topologized by the seminorms
@@
For $q=1$, as noted in Lemma 2.1 of [7] the connection to the trace norm is given by Theorem 9.3 of [1] and (C.3.1) of [10],
(7.1) $\|\mathrm{Op}_{h}^{w}(a(t))\|_{\mathcal{L}^{1}}\leq C_{n}\|a\|_{2n+1,S^{L^{1}}_{1/2}}$
Theorem 7.1.
Assume that $a\in S^{L^{1}}_{1/2}$, and let $a(t)=e^{tQ}a$. If $A(t)$ satisfies $\partial_{t}A(t)=\mathcal{L}A(t)$ with $A(0)=\mathrm{Op}_{h}^{w}(a)$, there is $C$ depending on $\|a_{0}\|_{N,S^{L^{1}}_{1/2}}$ for some NN, such that
$\|A(t)-\mathrm{Op}_{h}^{w}(a(t))\|_{\mathcal{L}^{1}}\leq Cth^{\frac{1}{2}},\quad 0\leq t<\infty.$
Proof.
The proof follows that of Theorem 2 of [7]. We include the outline of the proof for convenience. Using $A(0)=\mathrm{Op}_{h}^{w}(a(0))$ one writes
$A(t)-\mathrm{Op}_{h}^{w}(a(t))$ $=\int_{0}^{t}e^{(t-s)\mathcal{L}}(\mathcal{L}-\partial_{s})\mathrm{Op}_{h}^{w}(a(s))\,ds$
$=\int_{0}^{t}e^{(t-s)\mathcal{L}}\Bigl(\mathcal{L}\mathrm{Op}_{h}^{w}(a(s))-\mathrm{Op}_{h}^{w}(Qa(s))\Bigr)\,ds.$
@@
$\Bigl\|\mathcal{L}\mathrm{Op}_{h}^{w}(a(s))-\mathrm{Op}_{h}^{w}(Qa(s))\Bigr\|_{\mathcal{L}^{1}}\leq Ch^{\frac{1}{2}}.$
@@
$\sup_{s\geq 0}\|a(s)\|_{N,S_{1/2}^{L^{1}}}\leq C_{N}\,\|a\|_{N,S_{1/2}^{L^{1}}},\quad\forall N\geq 0.$
The jump operators $L_{j}$ are linear functions of $x$, so the composition rules for the semi-classical Weyl calculus imply that
(7.2) $\mathcal{L}\mathrm{Op}_{h}^{w}(a(s))-\mathrm{Op}_{h}^{w}(Qa(s))=\frac{i}{h}[P,\mathrm{Op}_{h}^{w}(a(s))]-\mathrm{Op}_{h}^{w}(H_{p}a(s)).$
@@
$\Bigl\|\frac{i}{h}[P,\mathrm{Op}_{h}^{w}(a(s))]-\mathrm{Op}_{h}^{w}(H_{p}a(s))\Bigr\|_{2n+1,S_{1/2}^{L^{1}}}\leq Ch^{\frac{1}{2}}\|a(s)\|_{N,S_{1/2}^{L^{1}}},$
together with (7.1). ∎
We now turn to results of [3], which concern bounds on the Hilbert-Schmidt norm of the difference, and relate to our estimates for $q=2$. The connection of $S_{\rho}^{L^{2}}$ to the Hilbert-Schmidt norm is given by
(7.3) $\|\mathrm{Op}_{h}^{w}(a(t))\|_{\mathcal{L}^{2}}=(2\pi h)^{-\frac{n}{2}}\|a(t)\|_{L^{2}}=(2\pi)^{-\frac{n}{2}}\|a(t)\|_{0,S^{L^{2}}_{\rho}}.$
In the following theorem the coupling constant $\gamma$ is allowed to depend on $h$, subject to lower bounds of $h^{\frac{1}{3}}$.
Theorem 7.2.
Assume that $\frac{1}{2}\leq\rho\leq\frac{2}{3}$, $a\in S^{L^{2}}_{\rho}$, and let $a(t)=e^{tQ}a$. Assume also that $\gamma=\gamma(h)$ satisfies $h^{2\rho-1}\leq\gamma\leq 1$. Then if $A(t)$ satisfies $\partial_{t}A(t)=\mathcal{L}A(t)$ with $A(0)=\mathrm{Op}_{h}^{w}(a)$, there is $C$ depending on $\|a_{0}\|_{N,S^{L^{2}}_{\rho}}$ for some NN, such that
$\|A(t)-\mathrm{Op}_{h}^{w}(a(t))\|_{\mathcal{L}^{2}}\leq Cth^{2-3\rho},\quad 0\leq t<\infty.$
Proof.
@@
$\Bigl\|\mathcal{L}\mathrm{Op}_{h}^{w}(a(s))-\mathrm{Op}_{h}^{w}(Qa(s))\Bigr\|_{\mathcal{L}^{2}}\leq Ch^{2-3\rho}.$
@@
$\Bigl\|\frac{i}{h}[P,\mathrm{Op}_{h}^{w}(a(s))]-\mathrm{Op}_{h}^{w}(H_{p}a(s))\Bigr\|_{0,S_{\rho}^{L^{2}}}\leq Ch^{2-3\rho}\|a(s)\|_{N,S_{\rho}^{L^{2}}}$

### v2

@@ 2. The model operator and the case $\epsilon=1$
is subelliptic since it satisfies the Hörmander condition, see [10] and [13]. By a result of Folland [4], $P_{0}(y,\partial_{y})$ has a unique fundamental solution that is homogeneous of degree 2-2, which in this case takes a simple form
(2.1) $K_{0}(y)=G_{y_{0}}(y^{\prime})G_{\frac{1}{12}y_{0}^{3}}(y^{\prime\prime}),$
@@ 3. The Fokker-Planck equation
Since $A_{2}(0,y^{\prime},y^{\prime\prime})=\partial_{y_{0}}A_{2}(0,y^{\prime},y^{\prime\prime})=0$, we can write $A_{2}(y)=y_{0}^{2}A_{0}(y)$, and similarly $A_{1}(y)=y_{0}A_{0}(y)$, in each case with $A_{0}\in C^{0}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},w}$.
@@ 5. Operator bounds
6. Proof of the subelliptic estimate
We establish (1.10) as a result of Theorem (6.1) below, which counts the order of derivatives in $x$ and $\xi$ as they are counted in the subelliptic calculus.
Let us first verify that (1.10) holds over $0\leq t\leq 1$. For $N=0$ this follows from the fact that $e^{tQ}$ is a contraction on $L^{p}$ for $1\leq p\leq\infty$ (see (6.3) below). Since the second order terms in $Q$ are constant coefficient, and all derivatives of $V$ and $\ell_{j}(x)$ are globally bounded, it follows that $Q,(\epsilon\partial_{x})^{\alpha}(\epsilon\partial_{\xi})^{\beta}]$ is a differential operator in $\epsilon\partial_{x}$ and $\epsilon\partial_{\xi}$ of order at most $|\alpha|+|\beta|$, with smooth bounded coefficients. An induction argument using the Duhamel formula and Gronwall’s inequality then establishes (1.10) uniformly over $t\in[0,1]$.
Combining this with (6.2) below for $T=1$ now shows that (1.10) holds globally in time.
Theorem 6.1.
Assume that $p=\frac{1}{2}|\xi|^{2}+V(x)$ with real $V$ satisfying (1.7), and $Q$ is of the form (1.8), where (1.9) holds with non-singular $B$. Then for all $N\in\mathbb{Z}_{+}$ there is $C_{N}$ such that, with $a(t)=e^{tQ}a(0)$ and $\epsilon=\sqrt{\gamma h/2}$, for all $1\leq p\leq\infty$, and all $\epsilon,\gamma\in(0,1]$,
(6.1) $\sup_{0\leq t\leq\infty}\,\sum_{|\alpha|+3|\beta|\leq 3N}\|(\epsilon\partial_{\xi})^{\alpha}(\epsilon\partial_{x})^{\beta}a(t)\|_{L^{p}(\mathbb{R}^{n})}\\ \leq C_{N}\sum_{|\alpha|+3|\beta|\leq 3N}\|(\epsilon\partial_{\xi})^{\alpha}(\epsilon\partial_{x})^{\beta}a(0)\|_{L^{p}(\mathbb{R}^{n})}.$
Additionally, for all $N\in\mathbb{Z}_{+}$ and $T>0$ there is $C_{N,T}$ so that
(6.2) $\sup_{T\leq t\leq\infty}\,\sum_{|\alpha|+3|\beta|\leq N}\|(\epsilon\partial_{\xi})^{\alpha}(\epsilon\partial_{x})^{\beta}a(t)\|_{L^{p}(\mathbb{R}^{n})}\leq C_{N,T}\|a(0)\|_{L^{p}(\mathbb{R}^{n})},$
where $C_{N,T}=\mathcal{O}(T^{-\frac{1}{2}N})$ as $T\rightarrow 0^{+}$.
For simplicity, in proving Theorem (6.1) we consider data $a(0)\in\mathcal{S}(\mathbb{R}^{2n})$; the a priori estimates will allow the extension of Theorem 6.1 to spaces of functions with $L^{p}$ bounds on a finite number of derivatatives. We start with the following two results.
Maximum Principal: $\sup_{t>0}\sup_{x,\xi}a(t,x,\xi)\leq\sup_{x,\xi}a(0,x,\xi).$
Mass Conservation: for $t>0$, $\;\int_{\mathbb{R}^{2n}}a(t,x,\xi)\,dx\,d\xi=\int_{\mathbb{R}^{2n}}a(0,x,\xi)\,dx\,d\xi.$
Both of these follow as for the standard heat equation; for completeness we include the proofs here. For the maximum principle, it suffices to work on a strip $t\in[0,T]$ for $T<\infty$. Given $c>0$ consider the function $a_{c}=a-ct$, so $(\partial_{t}-Q)a_{c}=-c<0$. At a local maximum of $a_{c}$ over $0\leq t\leq T$ we must have $\nabla_{x}a=0=\nabla_{\xi}a$, and $X_{j}^{2}a\leq 0$ for $j\geq 1$. Additionally $\partial_{t}a\geq 0$ if $t\neq 0$. This leads to a contradiction if $t\neq 0$. Thus $a_{c}$ attains it maximum at $t=0$. Since this holds for all $c>0$, it follows that $a$ attains its maximum at $t=0$.
@@ 6. Proof of the subelliptic estimate
$\partial_{t}\int a\,dx\,d\xi=\int Qa\,dx\,d\xi.$
The first order terms in $Q$ are divergence free, and the second order terms have constant coefficients, hence we may integrate by parts, using decay of $a$ in $(x,\xi)$, to see that the right hand side vanishes.
@@
$\|e^{tQ}a(0)\|_{L^{\infty}(\mathbb{R}^{2n})}\leq\|a(0)\|_{L^{\infty}(\mathbb{R}^{2n})}.$
@@
(6.3) $\|e^{tQ}a(0)\|_{L^{p}(\mathbb{R}^{2n})}\leq\|a(0)\|_{L^{p}(\mathbb{R}^{2n})}\quad\text{for all}\;\;1\leq p\leq\infty.$
@@
Given $N$, let $T_{K}$ be a right inverse for $\partial_{t}-Q$ as in (3.5) such that the remainder is of order 3N-3N, and let $g=T_{K}(\delta\otimes a(0))$. Since $K(v,y)-K_{0,\epsilon}(y)$ is of negative order, and $K_{0,\epsilon}(y_{0},\cdot)\rightarrow\delta(\cdot)$ as $y_{0}\rightarrow 0+$, by (5.1) we have that $\lim_{t\rightarrow 0+}g(t,\cdot)=a(0)$ in the $L^{p}(\mathbb{R}^{2n})$ norm.
@@
(6.4) $\sup_{t>0}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3m}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}g(t,\cdot)\bigr\|_{L^{p}}\\ \leq C_{N}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3m}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}a(0)\bigr\|_{L^{p}}.$
Additionally, $(\partial_{t}-Q)g=R_{3N+1}(\delta\otimes a(0))$ satisfies
(6.5) $\sup_{t>0}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3N}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}(\partial_{t}-Q)g(t,\cdot)\bigr\|_{L^{p}}\leq C_{N}\|a(0)\|_{L^{p}}.$
Let $h=e^{tQ}a(0)-g$, so $(\partial_{t}-Q)h=-(\partial_{t}-Q)g$ and $h=0$ for $t\leq 0$. To control $h$, let $K$ be the left parametrix constructed in Theorem 4.2 for $\partial_{t}-Q$ with remainder $R_{3N+1}$ of order 3N-3N, and write
@@
By (6.3) and (6.4) with $m=0$ we have $\|h(t,\cdot)\|_{L^{p}}\leq C\|a(0)\|_{L^{p}}$ for $t\geq 0$. If $|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq N$, and $w=(t,x,\xi)$, Theorem 5.5 with $m=0$ then yields
$\sup_{t>0}\,\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}\int R_{3N+1}(\epsilon,w,-y)h(\exp_{w}(y))\,dy\,\bigr\|_{L^{p}(dx\,d\xi)}\\ \leq C\|a(0)\|_{L^{p}},$
@@
$\sup_{t>0}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3N}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}h(t,\cdot)\bigr\|_{L^{p}(dx\,d\xi)}\leq C\|a(0)\|_{L^{p}}.$
@@ 7. Applications to classical-quantum correspondence
We show here how Theorem 1.2 and Theorem 1.3 follow from Theorem 6.1 and results from [11] and [5]. We consider real valued $V$ satisfying (1.7), and complex valued $\ell_{j}$ satisfying (1.9) with non-degenerate $B$, and define $Q$ by (1.8). Recall that $S^{L^{q}}_{\rho}$ is defined by the seminorms
@@
We first consider Theorem 1.2, which concerns $q=1$ and the trace-class norm. As noted in Lemma 2.1 of [11], the connection to the trace norm is given by Theorem 9.3 of [2] and (C.3.1) of [16],
(7.1) $\|\mathrm{Op}_{h}^{\rm w}(a)\|_{\mathcal{L}^{1}}\leq C_{n}\|a\|_{2n+1,S^{L^{1}}_{1/2}}$
Proof of Theorem 1.2.
The proof follows that of Theorem 2 of [11]. We include the outline of the proof for convenience. Using $A(0)=\mathrm{Op}_{h}^{\rm w}(a(0))$ one writes
$A(t)-\mathrm{Op}_{h}^{\rm w}(a(t))$ $=\int_{0}^{t}e^{(t-s)\mathcal{L}}(\mathcal{L}-\partial_{s})\mathrm{Op}_{h}^{\rm w}(a(s))\,ds$
$=\int_{0}^{t}e^{(t-s)\mathcal{L}}\Bigl(\mathcal{L}\mathrm{Op}_{h}^{\rm w}(a(s))-\mathrm{Op}_{h}^{\rm w}(Qa(s))\Bigr)\,ds.$
@@
$\Bigl\|\mathcal{L}\mathrm{Op}_{h}^{\rm w}(a(s))-\mathrm{Op}_{h}^{\rm w}(Qa(s))\Bigr\|_{\mathcal{L}^{1}}\leq Ch^{\frac{1}{2}}.$
@@
$\sup_{s\geq 0}\|a(s)\|_{N,S_{1/2}^{L^{1}}}\leq C_{N}\,\|a(0)\|_{N,S_{1/2}^{L^{1}}},\quad\forall N\geq 0.$
The jump operators $L_{j}$ are linear functions of $x$, so the composition rules for the semi-classical Weyl calculus imply that, with $P=p^{\rm w}(x,hD)$,
(7.2) $\mathcal{L}\mathrm{Op}_{h}^{\rm w}(a(s))-\mathrm{Op}_{h}^{\rm w}(Qa(s))=\frac{i}{h}[P,\mathrm{Op}_{h}^{\rm w}(a(s))]-\mathrm{Op}_{h}^{\rm w}(H_{p}a(s)).$
@@
$\Bigl\|\frac{i}{h}[P,\mathrm{Op}_{h}^{\rm w}(a(s))]-\mathrm{Op}_{h}^{\rm w}(H_{p}a(s))\Bigr\|_{2n+1,S_{1/2}^{L^{1}}}\leq Ch^{\frac{1}{2}}\|a(s)\|_{N,S_{1/2}^{L^{1}}},$
which together with (7.1) implies the desired estimate. ∎
Theorem 1.3 extends results of [5] concerning estimates in the Hilbert-Schmidt norm, and relate to our estimates for $q=2$. The connection of $S_{\rho}^{L^{2}}$ to the Hilbert-Schmidt norm is given by
(7.3) $\|\mathrm{Op}_{h}^{\rm w}(a(t))\|_{\mathcal{L}^{2}}=(2\pi h)^{-\frac{n}{2}}\|a(t)\|_{L^{2}}=(2\pi)^{-\frac{n}{2}}\|a(t)\|_{0,S^{L^{2}}_{\rho}}.$
Proof of Theorem 1.3.
@@
$\Bigl\|\mathcal{L}\mathrm{Op}_{h}^{\rm w}(a(s))-\mathrm{Op}_{h}^{\rm w}(Qa(s))\Bigr\|_{\mathcal{L}^{2}}\leq Ch^{2-3\rho}.$
@@
$\Bigl\|\frac{i}{h}[P,\mathrm{Op}_{h}^{\rm w}(a(s))]-\mathrm{Op}_{h}^{\rm w}(H_{p}a(s))\Bigr\|_{0,S_{\rho}^{L^{2}}}\leq Ch^{2-3\rho}\|a(s)\|_{N,S_{\rho}^{L^{2}}}$